OpenStudy (anonymous):

I am trying to figure out the slope of these two lines and if they are parallel. x+13=y y-x=-9

OpenStudy (anonymous):

solve the equations in terms of y first: y = x+13 second: y = x-9 since the equations are in the from y = mx+b, and m is the slope, they both have slopes of 1, which makes them parallel :D

OpenStudy (anonymous):

That is what I thought but I was not certain.. Thanks Daniel